# Nonequilibrium Phase Transitions in Large $N$ Matrix Quantum Mechanics

### v3

@@
Nonequilibrium Phase Transitions in
Large $N$ Matrix Quantum Mechanics
Minjae Cho
@@ 1 Introduction
Of course, whether the Lindblad equation provides a good effective description for an open quantum system depends on the specifics of the setup of interest. In particular, the Markovian property may not hold, and the environment may build a nontrivial memory of the subsystem’s evolution, in which case the Lindblad equation is not applicable. Finding or engineering specific setups in quantum gravity where the Lindblad equation provides a good effective description is beyond the scope of the current work. Instead, our primary interest here lies in exploring possible nonequilibrium phase transitions in quantum gravity, and the Lindblad equation offers an amenable setup for such investigations in the context of MQM.
@@
We now show that there is no Lindblad dissipation for planar gauged MQM. In the gauged MQM, well-defined linear operators are $U(N)$ singlets of the trace form ${\rm Tr\,}(\cdots)$, in which case the jump operators $L_{K}$ are also products of traces. Consider the time evolution of operator expectation values $\langle\cal O\rangle=\text{tr}\left(\rho{\cal O}\right)$, where tr denotes the trace over the Hilbert space, as governed by (1.1):
${d\over dt}\langle{\cal O}\rangle=i\langle[H,{\cal O}]\rangle+\sum_{K}\gamma_{K}\left(\langle L_{K}^{\dagger}{\cal O}L_{K}\rangle-{1\over 2}\langle L_{K}^{\dagger}L_{K}{\cal O}\rangle-{1\over 2}\langle{\cal O}L_{K}^{\dagger}L_{K}\rangle\right).$ (1.4)
In the planar limit $N\rightarrow\infty$, large $N$ factorization implies that the dissipation terms in (1.4) cancel exactly, and the dynamics reduces to unitary time evolution governed solely by the Hamiltonian. [^5] A potentially relevant statement regarding exact quantum error correcting codes in the context of planar gauged MQM was made in [15]. However, dissipative dynamics may become nontrivial at subleading orders in $1\over N$} or finite $N$.
Instead, we consider ungauged MQM, where $L_{K}$ are in non-singlet representations of $U(N)$. As a concrete example, define the adjoint annihilation and creation operators
@@
Now, the dissipation term is nontrivial in the planar limit and affects the dynamics. It can be regarded as a spontaneous emission of a single MQM quantum to a heat sink. This quantum is in the adjoint representation, sometimes referred to as the (off-shell) long strings [28, 29, 30, 31]. (1.7) respects the $U(N)$ symmetry in that the time evolution and the $U(N)$ transformations commute with each other:
@@
for all operators $\cal O$}. If steady states exist, then there must be at least one that transforms as a $U(N)$ singlet, i.e., $G_{ij},\rho]=0,~\forall i,j$. This does not necessarily imply that such a $\rho$ is an ensemble composed only of $U(N)$ singlet states in the Hilbert space, which would require $G_{ij}\rho=0$—a condition that fails in many examples we will consider. In the language of [32], density matrices satisfying $G_{ij}\rho=0$ are said to be strongly symmetric, while those satisfying $G_{ij},\rho]=0$ but not $G_{ij}\rho=0$ are weakly symmetric under $U(N)$. Suppose a steady state $\rho$ satisfies $G_{ij},\rho]=0$. If $\rho$ is pure, then it is strongly symmetric. Equivalently, if it is weakly but not strongly symmetric, then it must be a mixed state.
@@ 2.2 Quadratic case
The Hamiltonian (2.7) is nothing but that of $c=1$ and type 0B MQM. However, the vacua over which these theories are defined are quite different from the normalizable states of interest in this work. For the gauged model, the vacua of $c=1$ and type 0B MQM are given by filling the Fermi levels on one side and on both sides of the top of the potential, respectively, and the observables of interest are scattering amplitudes of fluctuations around the Fermi surface [50], which match the dual string theory amplitudes [51, 52, 53].
The adjoint sector of the ungauged model also allows for scattering states, called long strings, which match the open strings on FZZT branes in certain limits [29, 31]. In such discussions, long strings are regarded as fluctuations around the aforementioned singlet vacua. The investigation of the precise dual string theory description of the steady state at $\gamma>\gamma_{c}$ is beyond the scope of this work. Given that both $c=1$ and type 0B string theories in their current formulations describe perturbative scattering string states dual to fluctuations of the Fermi surface, it remains unclear whether there is a practical tool to describe the string background dual to the steady state.
@@ 3.1 Review of the bootstrap method
All bootstrap constraints are convex after the relaxation. We consider objective functions of the form $\langle{\rm Tr\,}{\cal Q}\rangle$, where ${\rm Tr\,}{\cal Q}$ is Hermitian and $\cal Q$} is a sum of words of lengths smaller than $L$. The final form of the bootstrap problem for steady states involving words of lengths up to $L$ is given by the following SDP problem:
@@ 7. Restrict to constraints involving only single- and double-traces in the above
When the strong symmetry constraints (3.9) are imposed, it may happen that the bootstrap problem SDP($L$) does not admit a feasible point, which can be detected by solvers such as MOSEK and SDPA-DD. In such cases, we conclude that no strongly symmetric steady state exists for the system under consideration, and in particular, that the steady state—if it exists—must be a mixed state.
@@ 5 Discussion
•  Even though there is no Lindblad dissipation for gauged MQM in the planar limit, it would be interesting to investigate the effects of dissipation either at finite $N$ or at subleading orders in the large $N$ expansion. The bootstrap approach can be straightforwardly extended to gauged MQM at finite $N$, where alternative methods based on suitable truncation schemes may also be available. A relevant remark is that, in the regime where a semiclassical Einstein gravity picture emerges, there is expected to be no essential difference between the gauged and ungauged MQM [30]. Away from such a regime, the two versions of MQM generally differ. The results presented in this work show that such differences are also present in the presence of Lindblad dissipative dynamics.

### v4

@@
It is believed that the theory of quantum gravity describing our universe is unitary. Nonetheless, if we only have access to a subsystem, its dynamics is described by nonequilibrium physics. Motivated by this, we investigate the planar limit of large $N$ ungauged one-matrix quantum mechanics obeying the Lindblad master equation with dissipative jump terms, focusing on the existence, uniqueness, and properties of steady states that signal nonequilibrium phase transitions. In the first class of examples, where potentials are unbounded from below, we study nonequilibrium critical points above which strong dissipation allows for the existence of normalizable steady states that would otherwise not exist. In the second class of examples, termed matrix quantum optics, we find evidence of nonequilibrium phase transitions analogous to those recently reported in the quantum optics literature for driven-dissipative Kerr resonators. Preliminary results on two-matrix quantum mechanics are also presented. We implement bootstrap methods to obtain concrete and rigorous results for the nonequilibrium steady states of matrix quantum mechanics in the planar limit.
@@ 1 Introduction
Of course, whether the Lindblad equation provides a good effective description for an open quantum system depends on the specifics of the setup of interest. In particular, the Markovian property may not hold, and the environment may build a nontrivial memory of the subsystem’s evolution, in which case the Lindblad equation is not applicable. Finding or engineering specific setups in quantum gravity where the Lindblad equation provides a good effective description is beyond the scope of the current work. In this work, we will assume that the Lindblad equation provides an appropriate description of the subsystem of interest. Our primary interest here lies in exploring possible nonequilibrium phase transitions in quantum gravity, and the Lindblad equation offers an amenable setup for such investigations in the context of MQM.
@@
We begin by considering the planar gauged MQM. In the gauged MQM, well-defined linear operators are $U(N)$ singlets of the trace form ${\rm Tr\,}(\cdots)$, in which case the jump operators $L_{K}$ are also products of traces. Consider the time evolution of operator expectation values $\langle\cal O\rangle=\text{tr}\left(\rho{\cal O}\right)$, where tr denotes the trace over the Hilbert space, as governed by (1.1):
${d\over dt}\langle{\cal O}\rangle$ $=i\langle[H,{\cal O}]\rangle+\sum_{K}\gamma_{K}\left(\langle L_{K}^{\dagger}{\cal O}L_{K}\rangle-{1\over 2}\langle L_{K}^{\dagger}L_{K}{\cal O}\rangle-{1\over 2}\langle{\cal O}L_{K}^{\dagger}L_{K}\rangle\right)$ (1.4)
$=i\langle[H,{\cal O}]\rangle+{1\over 2}\sum_{K}\gamma_{K}\left(\langle[L^{\dagger}_{K},{\cal O}]\rangle\langle L_{K}\rangle+\langle[{\cal O},L_{K}]\rangle\langle L^{\dagger}_{K}\rangle\right),$
where we used large $N$ factorization in the second line. We observe that, in general, there is nontrivial Lindblad dissipation for the planar gauged MQM, where the time evolution is nonlinear in the operator expectation values. [^5] We thank the anonymous referee of JHEP for pointing out that Lindblad dissipation may be nontrivial in the planar gauged MQM. For example, if $\mathcal{O}$ and $L_{K}$ are single-trace operators, ${d\over dt}\langle\mathcal{O}\rangle$ contains quadratic terms in the expectation values, even before taking into account any trace-cyclicity considerations to be discussed later (e.g. (3.4)). For this reason, analyzing the Lindblad dissipation of the planar gauged MQM is a highly involved task, which we will explore in a separate work [28].
In this work, we instead consider ungauged MQM, where $L_{K}$ are in non-singlet representations of $U(N)$. As a concrete example, define the adjoint annihilation and creation operators
@@
Now, the dissipation term is not manifestly nonlinear in the expectation values and is therefore more amenable to practical investigation compared to (1.4). [^6] Although (1.8) will involve nonlinear terms in the expectation values at the planar limit once trace cyclicity and operator algebra are taken into account (see section 3.1), its degree of non-convexity is nevertheless much milder than that of (1.4). This dissipation can be regarded as a spontaneous emission of a single MQM quantum to a heat sink. This quantum is in the adjoint representation, sometimes referred to as the (off-shell) long strings [29, 30, 31, 32]. (1.7) respects the $U(N)$ symmetry in that the time evolution and the $U(N)$ transformations commute with each other:
@@
for all operators $\cal O$}. If steady states exist, then there must be at least one that transforms as a $U(N)$ singlet, i.e., $G_{ij},\rho]=0,~\forall i,j$. This does not necessarily imply that such a $\rho$ is an ensemble composed only of $U(N)$ singlet states in the Hilbert space, which would further require $G_{ij}\rho=0$—a condition that fails in many examples we will consider. In the language of [33], density matrices satisfying $G_{ij}\rho=0$ are said to be strongly symmetric, while those satisfying $G_{ij},\rho]=0$ but not $G_{ij}\rho=0$ are weakly symmetric under $U(N)$. Suppose a steady state $\rho$ satisfies $G_{ij},\rho]=0$. If $\rho$ is pure, then it is strongly symmetric. Equivalently, if it is weakly but not strongly symmetric, then it must be a mixed state. Therefore, if there are only weakly symmetric steady states, we conclude that all steady states must be mixed states, a statement that has been useful in studying nonequilibrium phase transitions in recent literature (see e.g. [33, 4]).
@@ 2.2 Quadratic case
The Hamiltonian (2.7) is nothing but that of $c=1$ and type 0B MQM. However, the vacua over which these theories are defined are quite different from the normalizable states of interest in this work. For the gauged model, the vacua of $c=1$ and type 0B MQM are obtained in the double-scaling limit—rather than merely the ’t Hooft limit—by filling the Fermi levels on one side and on both sides of the top of the potential, respectively. The observables of interest are the scattering amplitudes of fluctuations around the Fermi surface [51], which match the dual string theory amplitudes [52, 53, 54].
The adjoint sector of the ungauged model also allows for scattering states, called long strings, which match the open strings on FZZT branes in certain limits [30, 32]. In such discussions, long strings are regarded as fluctuations around the aforementioned singlet vacua. The investigation of the precise dual string theory description of the steady state at $\gamma>\gamma_{c}$ is beyond the scope of this work. Given that both $c=1$ and type 0B string theories in their current formulations describe perturbative scattering string states dual to fluctuations of the Fermi surface in the double-scaling limit, it remains unclear whether there is a practical tool to describe the string background dual to the steady state. Firstly, there does not appear to be a natural double-scaling limit for the normalizable steady state, since there is no infinite Fermi sea to fill. Secondly, the normalizable steady state is obtained in the ’t Hooft limit, whereas perturbative string theory computations produce quantities only in the double-scaling limit. On another note, the non-singlet sector of MQM has been argued to be related to black holes in these string theories [30, 55, 56, 57]. It is therefore very interesting to study the Lindblad dynamics of MQM in the double-scaling limit, as opposed to the ’t Hooft limit considered in this work. We leave a detailed investigation of this question for future work.
@@ 3.1 Review of the bootstrap method
All bootstrap constraints are convex after the relaxation. We consider objective functions of the form $\langle{\rm Tr\,}{\cal Q}\rangle$, where ${\rm Tr\,}{\cal Q}$ is Hermitian and $\cal Q$} is a sum of words of lengths smaller than $L$. The final form of the bootstrap problem for steady states involving words of lengths up to $L$ is given by the following SDP problem: [^9] SDP is a convex optimization problem of finding the minimum (or maximum) of $\sum_{i=1}^{m}c_{i}x_{i}$ subject to the matrix positivity constraint $\textbf{X}=\sum_{i=1}^{m}F_{i}x_{i}-F_{0}\succeq 0$, where $x_{i}$ are real variables, $c_{i}$ are real constants, and $F_{0}$ and $F_{i}$ are real-symmetric constant matrices. For an introduction to SDP, we refer the reader to [61].
@@ 7. Restrict to constraints involving only single- and double-traces in the above
When the strong symmetry constraints (3.9) are imposed, it may happen that the bootstrap problem SDP($L$) does not admit a feasible point, which can be detected by solvers such as MOSEK and SDPA-DD. In such cases, we conclude that no strongly symmetric steady state exists for the system under consideration, and in particular, that the steady state—if it exists—must be a mixed state. [^11] To be precise, when SDP solvers produce infeasibility certificates, such as pdINF in SDPA-DD, this should be regarded only as numerical evidence for infeasibility, rather than a mathematically rigorous proof. For example, SDPA-DD treats $(x,\textbf{X},\textbf{Y})$ as variables, where Y is the dual variable appearing in the dual problem of maximizing ${\rm tr\,}(F_{0}\textbf{Y})$ subject to ${\rm tr\,}(F_{i}\textbf{Y})=c_{i}$ and $\textbf{Y}\succeq 0$, and always keeps the positivity constraints $\textbf{X}\succeq 0$ and $\textbf{Y}\succeq 0$ satisfied while attempting to minimize the residuals $\textbf{X}-\sum_{i=1}^{m}F_{i}x_{i}-F_{0}||$ and $\sum_{i=1}^{m}({\rm tr\,}(F_{i}\textbf{Y})-c_{i})^{2}$, which serve as indicators of infeasibility, while simultaneously reducing the duality gap via Newton-type methods. During this process, however, step lengths may be significantly reduced to maintain positivity constraints even when the residuals or duality gap remain relatively large, and such behavior is then interpreted by the solver as evidence of infeasibility. Another relevant point is that the scale of the initial starting point is set by SDPA parameters, such as lambdaStar, which in general should be taken large enough (see 6.2 of [66] for the precise definition of pdINF, which indicates infeasibility only within a subregion of $\textbf{X}\succeq 0$ and $\textbf{Y}\succeq 0$ determined by the SDPA parameters). To ensure that one does not miss a potentially very small feasible region in this procedure, higher-precision solvers are desirable for infeasibility tests, which is why we used SDPA-DD for such tests.
@@ 5 Discussion
•  Although more involved, the bootstrap approach can be straightforwardly extended to gauged MQM, where alternative methods based on suitable truncation schemes may also be available. A relevant remark is that, in the regime where a semiclassical Einstein gravity picture emerges, there is expected to be no essential difference between the gauged and ungauged MQM [31]. Away from such a regime, the two versions of MQM generally differ. The results presented in this work suggest that such differences may also be present in the presence of Lindblad dissipative dynamics.